# Large deviations in statistics of the local time and occupation time for a run and tumble particle

### v1

@@ I Introduction
While the local time and occupation time have been studied extensively for Brownian particles CDM2002 ; PCRK2019 ; BMR23 ; SmithMeerson24 , there are very few existing results in the biologically-relevant context of active matter Schweitzer2003 ; RBELS2012 ; MJRLPRS2013 ; FGGVWW2015 ; GKKRST23 . The study of fluctuations of local time for active particles is relevant in recent research as their self-propelled motion useful in molecular transport Santiago2018 and drug delivery GXGG2020 at the nanoscale. In Ref. SK21 the local time for a run and tumble (RTP) in one dimension was studied. They obtained an exact result for the distribution of the local time, in the absence of an external drift. They then extracted the behavior in the regime of typical fluctuations at long times. The large-deviation regime was not studied in detail.
@@
where $x(t)$ is a stochastic process (particular examples will be given below), and $f(x)$ is some function. In the limit $T\to\infty$, it follows from ergodicity that the time average and ensemble average coincide, and therefore, with probability one, $\mathcal{A}/T$ will equal its corresponding ensemble-average value. However, if $T$ is large but finite, there will be fluctuations away from this expected value which are of interest.
For finite observation time $T$ its probability density function (PDF), $P(\mathcal{A})$, depends on details such that the initial position $x_{0}=x(0)$ of the Brownian motion. However in the long $T$ limit, it satisfies a universal large deviation principle (LDP) Hugo2009
$P(\mathcal{A})\sim\,e^{-T\,I(\mathcal{A}/T)}$ (2)
@@
$I(\rho)=-\lim_{T\to\infty}\frac{1}{T}\ln P(\mathcal{A}=\rho T)$ (3)
The rate function $I(\rho)$ can be calculated via the Legendre-Fenchel transform of the scaled cumulant generating function (SCGF), which is defined as
$\lambda(k)=\lim_{T\to\infty}\frac{1}{T}\,\,\ln\langle e^{k\,\mathcal{A}}\rangle\,.$ (4)
@@
$I(\rho)=\sup_{k}\{k\,\rho-\lambda(k)\}$ (6)
provided that $\lambda(k)$ exists and is differentiable.
@@
In particular, the DV formalism was used successfully in Refs. NT2018 ; NT2018DPT to calculate large deviations of the occupation time of a drifted Brownian particle in a finite interval. Remarkably, they found that (for nonzero drift) the associated rate function contains a singularity. Such singularities are usually interpreted as dynamical phase transitions (DPT), in analogy with equilibrium statistical physics exclusion ; glass ; kafri ; exclusion1 ; singularities ; baek ; baek1 ; NT2018 ; NT2018DPT ; MukherjeeSmith23 ; Smith22Chaos ; MukherjeeSmithConvexhull .
The two observables under study in this work are particular cases of Eq. (1): local time and occupation time, defined as
@@
$\mathds{1}_{[-a,a]}(x)=\begin{cases}1&\text{for}-a\leq x\leq a\\[2.84526pt] 0&\text{otherwise}\\ \end{cases}$ (9)
The mean values of these observables scale, at long times, as $\sim\sqrt{T}$ in the absence of a drift, and as $\sim T^{0}$ if a drift is present. We study here the large deviations regime that describes fluctuations of order $\sim T$ of the local time and occupation time for RTP’s, with or without external drifts, i.e., we focus here on the right tails of the distributions of $\mathcal{T}$ and $\mathcal{R}$. In particular, the probability that $\mathcal{R}=T$ is the survival probability Weiss84 ; WMLW87 ; MPW92 ; MalakarEtAl18 ; MLMS20 ; BMS21 of the particle inside the interval. In fact, in doing so, we will also extend the general theory of large deviations of dynamical observables (1) (i.e., the DV theory) to the case in which $x(t)$ describes the motion of an RTP in one dimension. As in the case of the Brownian motion, here too we find that, in the case of nonzero drift, the rate functions describing the local time or occupation time distribution contain singularities which we interpret as first-order DPTs.
@@ II Local time for Brownian motion with drift
$\mathcal{L}_{k}=\mathcal{L}+k\;\delta(x)$ (13)
@@
$\mathcal{L}=D\;\partial_{x}^{2}+\mu\,\partial_{x}$ (14)
@@
$\mathcal{H}=D\,\partial_{x}^{2}-\frac{\mu^{2}}{4D}+k\,\delta(x)$ (16)
@@
$\mathcal{H}\,\psi(x)=\tilde{\lambda}(k,\mu)\,\psi(x)$ (17)
@@
$\tilde{\lambda}(k,\mu)=\tilde{\lambda}(k,0)-\frac{\mu^{2}}{4D}$ (19)
@@ III Local time for RTP with drift
$\mathcal{T}=\frac{1}{v_{0}+\mu}\,N_{+}(T)+\frac{1}{v_{0}-\mu}\,N_{-}(T)$ (24)
@@
Since $N(T)$ is bounded by the total number of tumbles $\mathcal{N}(T)$: $N(T)\leq\mathcal{N}(T)$, and as a result, $\left\langle e^{kN(T)}\right\rangle\leq\left\langle e^{k\mathcal{N}(T)}\right\rangle$. As we now show, this gives a upper bound for $\Lambda(k)$.
@@
$\langle e^{k\,\mathcal{N}(T)}\rangle=\sum_{\mathcal{N}=0}^{\infty}\,\frac{1}{\mathcal{N}!}\,e^{k\,\mathcal{N}}\,e^{-\gamma\,T}\,\left(\gamma\,T\right)^{\mathcal{N}}=e^{\gamma\,T\,\left(e^{k}-1\right)}\,.$ (28)
@@
$\tilde{I}\left(a\right)\geq a\ln\left(a/\gamma\right)-a+\gamma$ (32)
@@
$\partial_{t}\,\left(\begin{array}[]{c}P_{+}\\ P_{-}\end{array}\right)\,=\,L^{\dagger}\,\left(\begin{array}[]{c}P_{+}\\ P_{-}\end{array}\right)$ (33)
@@
$\mathcal{G}_{k\pm}(x,t)=\sum_{i=0}^{\infty}\,c_{i}\,e^{\xi_{i}\,t}\,\,r_{k\pm}^{(i)}(x)$ (43)
where $\xi_{i}$ and $r_{k}^{(i)}(x)$ are the eigenvalues and right eigenvectors of the generator respectively, $L^{\dagger}_{k}r_{k}=\xi\,r_{k}$, with the boundary conditions (42). In the long-$T$ limit the term with largest eigenvalue $\xi_{\max}$ dominates in the above sum (43). Using that
@@
In addition to the boundary conditions at $x=0$, the eigenfunctions must also satisfy boundary conditions at $x\to\pm\infty$. Eqs. (47) and (48) are linear, homogeneous, first-order ordinary differential equations with constant coefficients and may therefore be solved immediately in each of the two regimes $x>0$ and $x<0$. The solutions, that vanish at $x\to\pm\infty$, are given by
@@
$\alpha_{1,2}=\frac{1}{v_{0}^{2}-\mu^{2}}\,\left[\mu(\tilde{\Lambda}+\gamma)\pm\sqrt{\gamma^{2}\,\mu^{2}+\tilde{\Lambda}\,v_{0}^{2}\,(2\gamma+\tilde{\Lambda})}\right]\,.$ (53)
@@ III.2 Nonzero drift $\mu\neq 0$
For $\mu\neq 0$, as in the case of the SCGF that describes the local-time statistics for Brownian motion, $\tilde{\Lambda}(k)$ becomes negative below a critical value of $k=k_{c}$. Following a similar argument to the one used above in Section II and in NT2018 ; NT2018DPT , the correct SCGF must satisfy be convex, and satisfy $\Lambda(0)=0$ and $\Lambda^{\prime}(0)=0$. Since zero is also an eigenvalue of the tilted generator, one finds that $\Lambda\left(k\right)=\max\left\{0,\tilde{\Lambda}\left(k\right)\right\}$, i.e., (using Eq. (56)):
@@
$\tilde{\rho}=\tilde{\rho}_{c}=\frac{\gamma\,\mu}{v_{0}}$ (64)
@@
$k_{c}\simeq\frac{1}{2}\ln\left(\frac{2}{1-\frac{\mu}{v_{0}}}\right)\gg 1$ (68)
@@
Remark. For $\mu=0$ one can extract the large time $T$ behavior of the cumulants by taking derivatves of (57) at $k=0$. One finds that for even integer $q$ they behave as
$\langle N(T)^{q}\rangle_{c}\simeq\gamma T\quad,\quad\langle{\cal T}^{q}\rangle=\frac{\gamma T}{v_{0}^{q}}$ (70)
@@ IV.1 General
$\left(\begin{array}[]{c}\dot{G}_{k+}\\ \dot{G}_{k-}\end{array}\right)=\mathcal{L}_{k}^{\dagger}\left(\begin{array}[]{c}G_{k+}\\ G_{k-}\end{array}\right)$ (77)
@@
$\mathcal{Q}\left(\begin{array}[]{c}\Phi_{1}\\ \Phi_{2}\end{array}\right)=\lambda\left(\begin{array}[]{c}\Phi_{1}\\ \Phi_{2}\end{array}\right)$ (82)
@@ IV.2 Occupation time
$\qquad\qquad\qquad=\frac{\lambda(k)\left[\lambda(k)-k+2\gamma\right]}{\left[\lambda(k)+2\gamma\right]\left[k-\lambda(k)\right]}$ (90)
@@
$\qquad\qquad\qquad=\frac{\lambda(k)\left[k-\lambda\left(k\right)-2\gamma\right]}{\left[\lambda(k)+2\gamma\right]\left[k-\lambda(k)\right]}$ (91)
@@
$\lambda(k)=2\frac{a^{2}\gamma}{v_{0}^{2}}k^{2}+O(k^{3})$ (92)
$I(\rho)=\frac{v_{0}^{2}\rho^{2}}{8\gamma a^{2}}+O(\rho^{3})$ (93)
A few higher order terms, as well as up to the fourth cumulant of ${\cal R}$ are given in the Appendix B. For $\rho\to 1$ (corresponds to large $k\to\infty$) the behaviour of the rate function depends on $a$. For details see Appendix B.1. This $I(\rho=1)$ can also be interpreted as the long-time survival rate of the RTP inside the interval $[-a,a]$ (see Appendix C). It is given by the solution to the transcendental equation
$\frac{a\gamma}{v_{0}}=\frac{\cos^{-1}(-1+I(1))}{2\sqrt{I(1)(2-I(1))}}$ (94)
for $\tilde{a}>1/2$ (see analgous formula for $\tilde{a}<1/2$ in Appendix B.1).
Note that the rate functions depend on a single dimensionless parameter $\tilde{a}=a\gamma/v_{0}$. We now discuss separately the two limits $\tilde{a}\gg 1$ and $\tilde{a}\ll 1$
@@ IV.5 Distribution of the endpoint conditioned on $\rho$
Physically, $\Phi_{1}(x)$ (properly normalized) gives the distribution of the RTP’s position at time $T$, conditioned on a given value of $\mathcal{R}$, cf. CH15 . From Eq. (84) one finds that if $\alpha^{2}>0$ then this distribution is unimodal and maximum at the origin, whereas if $\alpha^{2}<0$ it is bimodal and maximum at the edges of the interval. As we will see below, this change of behavior occurs only when $\tilde{a}=a\gamma/v_{0}<1/2$ and in that case for $\rho=\rho_{c}(\tilde{a})$ given by
$\rho_{c}(\tilde{a})=\frac{8\tilde{a}^{2}\left(3-4\tilde{a}^{2}\right)}{16\tilde{a}^{4}+3}\quad$ (98)
@@ V Summary and Discussion
Several future directions of research remain. It would be interesting to further understand the large-deviations behavior for the local time and occupation time for RTP by characterizing the conditioned process, i.e., the dynamics of the RTP conditioned on a given value of local time or occupation time, in greater detail than in the present work. It would also be interesting to study other dynamical observables of the type (1), for RTPs that are perhaps trapped in external potentials. One could also study dynamical observables that are of a more general form than the type (1) Hugo2018 or even study more detailed characterizations of the process such as the empirical position distribution (so-called level 2 large deviations HNE16 ; CVC22 ; Monthus24 ). The DV formalism can be extended for other types of active particles, and perhaps in higher dimension (e.g., active Brownian particles BLLRVV2016 ). It would also be interesting to study multi-particle systems, with possible interactions between the particles LMS19 ; LMS21 ; Singh21 ; ARYL21 ; PBDN21 ; RSBI22 ; Cates22 ; ARYL22 ; KPS23 ; MukherjeeSmith23 ; SmithMeerson24 .

### v2

@@ I Introduction
While the local time and occupation time have been studied extensively for Brownian particles CDM2002 ; PCRK2019 ; BMR23 ; SmithMeerson24 , there are very few existing results in the biologically-relevant context of active matter Schweitzer2003 ; RBELS2012 ; MJRLPRS2013 ; FGGVWW2015 ; GKKRST23 . The time spent by drug molecules in the vicinity of their target is of crucial importance for their efficiency, and it has been argued that the activity of drug molecules may enhance their ability to reach the target Santiago2018 ; GXGG2020 . In Ref. SK21 the local time for a run and tumble particle (RTP) in one dimension was studied. They obtained an exact result for the distribution of the local time, in the absence of an external drift. They then extracted the behavior in the regime of typical fluctuations at long times. The large-deviation regime was not studied in detail.
@@
where $x(t)$ is a stochastic process (particular examples will be given below), and $f(x)$ is some function. In the limit $T\to\infty$, assuming the process is ergodic, the time average and ensemble average coincide, and therefore, with probability one, $\mathcal{A}/T$ will equal its corresponding ensemble-average value. However, if $T$ is large but finite, there will be fluctuations away from this expected value which are of interest.
For finite observation time $T$ its probability density function (PDF), $P(\mathcal{A})$, depends on details such as the initial position $x_{0}=x(0)$ of the Brownian motion. However in the long $T$ limit, it satisfies a universal large deviation principle (LDP) Hugo2009
$P(\mathcal{A})\sim\,e^{-T\,I(\mathcal{A}/T)},$ (2)
@@
$I(\rho)=-\lim_{T\to\infty}\frac{1}{T}\ln P(\mathcal{A}=\rho T).$ (3)
According to Gärtner-Ellis theorem, the rate function $I(\rho)$ can be calculated via the Legendre-Fenchel transform of the scaled cumulant generating function (SCGF) Hugo2009 , which is defined as
$\lambda(k)=\lim_{T\to\infty}\frac{1}{T}\,\,\ln\langle e^{k\,\mathcal{A}}\rangle\,,$ (4)
@@
$I(\rho)=\sup_{k}\{k\,\rho-\lambda(k)\},$ (6)
provided that $\lambda(k)$ exists and is differentiable. If $\lambda(k)$ is not differentiable, then $I(\rho)$ is replaced by its convex envelope Hugo2009 .
@@
In particular, the DV formalism was used successfully in Refs. NT2018 ; NT2018DPT to calculate large deviations of the occupation time of a drifted Brownian particle in a finite interval. Remarkably, they found that (for nonzero drift) the associated rate function contains a singularity. Such singularities are usually interpreted as dynamical phase transitions (DPT), in analogy with equilibrium statistical physics exclusion ; glass ; kafri ; exclusion1 ; singularities ; baek ; baek1 ; NT2018 ; NT2018DPT ; MukherjeeSmith23 ; Smith22Chaos ; MukherjeeSmithConvexhull . Large deviations of the local time for a discrete model of a RTP on a lattice were studied in MBE19 . In the presence of an external drift, a DPT was uncovered, which is qualitatively similar to the DPT from the Brownian case NT2018 ; NT2018DPT .
In this paper we consider a continuum model for a RTP in one dimension. The two observables under study in this work are particular cases of Eq. (1): local time and occupation time, defined as
@@
$\mathds{1}_{[-a,a]}(x)=\begin{cases}1&\text{for }-a\leq x\leq a,\\[2.84526pt] 0&\text{otherwise}.\\ \end{cases}$ (9)
The mean values of these observables scale, at long times, as $\sim\sqrt{T}$ in the absence of a drift, and as $\sim T^{0}$ if a drift is present. We study here the large deviations regime that describes fluctuations of order $\sim T$ of the local time and occupation time for RTP’s, with or without external drifts, i.e., we focus here on the right tails of the distributions of $\mathcal{T}$ and $\mathcal{R}$. In particular, the probability that $\mathcal{R}=T$ is the survival probability Weiss84 ; WMLW87 ; MPW92 ; MalakarEtAl18 ; MLMS20 ; BMS21 ; NS24 of the particle inside the interval. In fact, in doing so, we will also extend the general theory of large deviations of dynamical observables (1) (i.e., the DV theory) to the case in which $x(t)$ describes the motion of an RTP in one dimension. As in the case of the Brownian motion, here too we find that, in the case of nonzero drift, the rate functions describing the local time or occupation time distribution contain singularities which we interpret as first-order DPTs.
@@ II Local time for Brownian motion with drift
$\mathcal{L}_{k}=\mathcal{L}+k\;\delta(x),$ (13)
@@
$\mathcal{L}=D\;\partial_{x}^{2}+\mu\,\partial_{x},$ (14)
@@
$\mathcal{H}=D\,\partial_{x}^{2}-\frac{\mu^{2}}{4D}+k\,\delta(x),$ (16)
@@
$\mathcal{H}\,\psi(x)=\tilde{\lambda}(k,\mu)\,\psi(x),$ (17)
@@
$\tilde{\lambda}(k,\mu)=\tilde{\lambda}(k,0)-\frac{\mu^{2}}{4D},$ (19)
@@ III Local time for RTP with drift
$\mathcal{T}=\frac{1}{v_{0}+\mu}\,N_{+}(T)+\frac{1}{v_{0}-\mu}\,N_{-}(T),$ (24)
@@
Since $N(T)$ is bounded by the total number of tumbles $\mathcal{N}(T)$ (i.e $N(T)\leq\mathcal{N}(T)$), it follows that $\left\langle e^{kN(T)}\right\rangle\leq\left\langle e^{k\mathcal{N}(T)}\right\rangle$ (for $k\geq 0$). As we now show, this gives a upper bound for $\Lambda(k)$.
@@
$\langle e^{k\,\mathcal{N}(T)}\rangle=\sum_{\mathcal{N}=0}^{\infty}\,\frac{1}{\mathcal{N}!}\,e^{k\,\mathcal{N}}\,e^{-\gamma\,T}\,\left(\gamma\,T\right)^{\mathcal{N}}=e^{\gamma\,T\,\left(e^{k}-1\right)}\,,$ (28)
@@
$\tilde{I}\left(a\right)\geq a\ln\left(a/\gamma\right)-a+\gamma,$ (32)
@@
$\partial_{t}\,\left(\begin{array}[]{c}P_{+}\\ P_{-}\end{array}\right)\,=\,L^{\dagger}\,\left(\begin{array}[]{c}P_{+}\\ P_{-}\end{array}\right),$ (33)
@@
$\mathcal{G}_{k\pm}(x,t)=\sum_{i=0}^{\infty}\,c_{i}\,e^{\xi_{i}\,t}\,\,r_{k\pm}^{(i)}(x),$ (43)
where $\xi_{i}$ and $r_{k}^{(i)}(x)$ are the eigenvalues and right eigenvectors of the generator respectively, $L^{\dagger}_{k}r_{k}=\xi\,r_{k}$, with the boundary conditions (42). Note that if the spectrum of $L^{\dagger}_{k}$ is continuous, then the sum (43) becomes an integral. In the long-$T$ limit the term with largest eigenvalue $\xi_{\max}$ dominates in the above sum (43). Using that
@@
In addition to the boundary conditions at $x=0$, the eigenfunctions corresponding to the largest eigenvalue must also satisfy boundary conditions at $x\to\pm\infty$. Eqs. (47) and (48) are linear, homogeneous, first-order ordinary differential equations with constant coefficients and may therefore be solved immediately in each of the two regimes $x>0$ and $x<0$. The solutions, that vanish at $x\to\pm\infty$, are given by
@@
$\alpha_{1,2}=\frac{1}{v_{0}^{2}-\mu^{2}}\,\left[\mu(\tilde{\Lambda}+\gamma)\pm\sqrt{\gamma^{2}\,\mu^{2}+\tilde{\Lambda}\,v_{0}^{2}\,(2\gamma+\tilde{\Lambda})}\right]\,,$ (53)
@@ III.2 Nonzero drift $\mu\neq 0$
For $\mu\neq 0$, as in the case of the SCGF that describes the local-time statistics for Brownian motion, $\tilde{\Lambda}(k)$ becomes negative below a critical value of $k=k_{c}$. Following a similar argument to the one used above in Section II and in NT2018 ; NT2018DPT , the correct SCGF must be convex, and satisfy $\Lambda(0)=0$ and $\Lambda^{\prime}(0)=0$. Since zero is also an eigenvalue of the tilted generator (see Ref. NT2018DPT and Appendix E), one finds that $\Lambda\left(k\right)=\max\left\{0,\tilde{\Lambda}\left(k\right)\right\}$, i.e., (using Eq. (56)):
@@
$\tilde{\rho}=\tilde{\rho}_{c}=\frac{\gamma\,\mu}{v_{0}},$ (64)
@@
$k_{c}\simeq\frac{1}{2}\ln\left(\frac{2}{1-\frac{\mu}{v_{0}}}\right)\gg 1,$ (68)
@@
Remark. For $\mu=0$ one can extract the large time $T$ behavior of the cumulants by taking derivatives of (57) at $k=0$. One finds that for even integer $q$ they behave as
$\langle N(T)^{q}\rangle_{c}\simeq\gamma T\quad,\quad\langle{\cal T}^{q}\rangle=\frac{\gamma T}{v_{0}^{q}}.$ (70)
@@ IV.1 General
$\left(\begin{array}[]{c}\dot{G}_{k+}\\ \dot{G}_{k-}\end{array}\right)=\mathcal{L}_{k}^{\dagger}\left(\begin{array}[]{c}G_{k+}\\ G_{k-}\end{array}\right),$ (77)
@@
$\mathcal{Q}\left(\begin{array}[]{c}\Phi_{1}\\ \Phi_{2}\end{array}\right)=\lambda\left(\begin{array}[]{c}\Phi_{1}\\ \Phi_{2}\end{array}\right),$ (82)
@@ IV.2 Occupation time
$\qquad\qquad\qquad=\frac{\lambda(k)\left[\lambda(k)-k+2\gamma\right]}{\left[\lambda(k)+2\gamma\right]\left[k-\lambda(k)\right]},$ (90)
@@
$\qquad\qquad\qquad=\frac{\lambda(k)\left[k-\lambda\left(k\right)-2\gamma\right]}{\left[\lambda(k)+2\gamma\right]\left[k-\lambda(k)\right]},$ (91)
@@
$\lambda(k)=2\frac{a^{2}\gamma}{v_{0}^{2}}k^{2}+O(k^{3}),$ (92)
$I(\rho)=\frac{v_{0}^{2}\rho^{2}}{8\gamma a^{2}}+O(\rho^{3}).$ (93)
A few higher order terms, as well as up to the fourth cumulant of occupation time ${\cal R}$ (8) are given in the Appendix B. For $\rho\to 1$ (corresponding to large $k\to\infty$) the behaviour of the rate function depends on $a$. For details see Appendix B.1. This $I(\rho=1)$ can also be interpreted as the long-time survival rate of the RTP inside the interval $[-a,a]$ (see Appendix C). It is given by the solution to the transcendental equation
$\frac{a\gamma}{v_{0}}=\frac{\cos^{-1}(-1+I(1))}{2\sqrt{I(1)(2-I(1))}}.$ (94)
for $\tilde{a}>1/2$, where $\tilde{a}=a\gamma/v_{0}$ (see analogous formula for $\tilde{a}<1/2$ in Appendix B.1).
Note that the rate functions depend on the single dimensionless parameter $\tilde{a}$. We now discuss separately the two limits $\tilde{a}\gg 1$ and $\tilde{a}\ll 1$
@@ IV.5 Distribution of the endpoint conditioned on $\rho$
Physically, $\Phi_{1}(x)$ (properly normalized) gives the distribution of the RTP’s position at time $T$, conditioned on a given value of $\mathcal{R}$, cf. CH15 . From Eq. (84) one finds that if $\alpha^{2}>0$ then this distribution is unimodal and maximal at the origin, whereas if $\alpha^{2}<0$ it is bimodal and maximal at the edges of the interval. As we will see below, this change of behavior occurs only when $\tilde{a}=a\gamma/v_{0}<1/2$ and in that case for $\rho=\rho_{c}(\tilde{a})$ given by
$\rho_{c}(\tilde{a})=\frac{8\tilde{a}^{2}\left(3-4\tilde{a}^{2}\right)}{16\tilde{a}^{4}+3},\quad$ (98)
@@ V Summary and Discussion
Several future directions of research remain. It would be interesting to further understand the large-deviations behavior for the local time and occupation time for RTP by characterizing the conditioned process, i.e., the dynamics of the RTP conditioned on a given value of local time or occupation time, in greater detail than in the present work MM22 ; MBE19 . It would also be interesting to study other dynamical observables of the type (1), for RTPs that are perhaps trapped in external potentials. One could also study dynamical observables that are of a more general form than the type (1) Hugo2018 or even study more detailed characterizations of the process such as the empirical position distribution (so-called level 2 large deviations HNE16 ; CVC22 ; Monthus24 ). The DV formalism can be extended for other types of active particles, and perhaps in higher dimension (e.g., active Brownian particles BLLRVV2016 ). It would also be interesting to study multi-particle systems, with possible interactions between the particles LMS19 ; LMS21 ; Singh21 ; ARYL21 ; PBDN21 ; RSBI22 ; Cates22 ; ARYL22 ; KPS23 ; MukherjeeSmith23 ; SmithMeerson24 .